# Towards understanding evolution of science through language model series

### v1

@@ Abstract
We introduce AnnualBERT, a series of language models designed specifically to capture the temporal evolution of scientific text. Deviating from the prevailing paradigms of subword tokenizations and “one model to rule them all”, AnnualBERT adopts whole words as tokens and is composed of a base RoBERTa model pretrained from scratch on the full-text of 1.7 million arXiv papers published until 2008 and a collection of progressively trained models on arXiv papers at an annual basis. We demonstrate the effectiveness of AnnualBERT models by showing that they not only have comparable performances in standard tasks but also achieve state-of-the-art performances on domain-specific NLP tasks as well as link prediction tasks in the arXiv citation network. We then utilize probing tasks to quantify the models’ behavior in terms of representation learning and forgetting as time progresses. Our approach enables the pretrained models to not only improve performances on scientific text processing tasks but also to provide insights into the development of scientific discourse over time. The series of the models is available at https://huggingface.co/jd445/AnnualBERTs.
I Introduction
@@
Natural language processing (NLP) has been experiencing a new paradigm where Transformers-based language models pretrained on a massive text corpus have significantly improved performances in downstream NLP tasks [1]. In our focused scientific domains, both the encoder and decoder component of Transformers have been extensively employed to develop scientific language models pretrained on full-text of papers [2].
@@
When developing language models for a target domain, prior studies have suggested a number of pretraining methods. The first is continual training, which uses a model pretrained already on a general domain corpus and continues the pretraining process on the target domain corpus (i.e., domain adaption, mixed-domain pretraining) [3]. BioBERT is one example adapted from the original BERT to the biomedical domain using the PubMed corpus [4]. The second approach is training from scratch directly on the target domain corpus. Recent development of BERT-like models, including both multi-domain [5, 6] and domain-specific ones [7, 8, 9], adopts this strategy, each aimed to better capture linguistic and semantic features of their respective fields. Training from scratch has been shown to over-perform continual training, and one reason for the over-performance is because it constructs a vocabulary encompassing more in-domain tokens, leading to more effective learning from the domain corpus. Continual training, by contrast, uses the same vocabulary from the original model and consequently fragments complex, domain-relevant words into less meaningful pieces. For example, “chloramphenicol” is treated as one token by BiomedBERT (PubmedBERT) but six tokens by BioBERT (“ch-lor-amp-hen-ico-l”).
@@
Here we propose a novel pretraining strategy that consists of two stages: first domain specific pretraining from scratch followed by progressively continual training on corpora within the same domain organized by year (Fig. 1). Our method is designed to, in addition to recognizing the importance of domain specific pretraining, capture the temporal evolution of language use. Temporal change is one of the most salient features in real-world language data: Scientific papers are published in a chronological order, which collectively represent the evolving landscape of scientific knowledge during its development; similarly, news articles and social media posts are produced over time, the collection of which shapes the global news agenda. However, existing efforts, regardless of the training methods, largely ignore the temporal aspect of language data and instead are directed towards developing “one model to rule them all”—one model that can excel in various downstream scientific NLP tasks like named entity recognition.
@@
Figure 1: Workflow for developing our AnnualBERT models pretrained on arXiv papers. We first train from scratch a RoBERTa model on papers published until 2008, denoted as $\mathcal{M}_{\text{base}}$, and then for each subsequent year $t$ from 2009, we use papers published in that year for continual training of $\mathcal{M}_{t-1}$ to obtain $\mathcal{M}_{t}$. During the training process, we use whole word token, represented as $\text{WT}_{1}$, $\text{WT}_{2}$, …, $\text{WT}_{n}$.
@@
By comparison, our proposal enables developing a series of language models on scientific documents, each of which represents one historical period and the collection of which allows us to understand the evolution of science. At a broader level, our vision is that a language model itself can be regarded as a highly knowledgeable researcher with extensive reading capabilities across all the scientific fields, who updates its inner workings (i.e., model parameters) upon reading all the papers published in a year. By probing how different versions of the model perform on the same task and by mining how model parameters change, we may be able to uncover how science evolves temporally. From another perspective, language models serve to compress documents over a certain period, thereby reflecting various characteristics of language use during that time, such as vocabulary, semantics, and topics. In this regard, our work is related to some recent studies that identify the temporal misalignment phenomenon [10, 11] and the remedies to derive customized models to different periods [11, 12, 13]. However, the language models used in the first place was pretrained on a time discorded corpus, which confounds the derivation.
@@
  • We evaluate AnnualBERT models on a suit of standard NLP tasks in the scientific domain, showing comparable performances to similar models. Moreover, we formulate several domain-specific tasks and demonstrate AnnualBERT has better performances for these in-domain tasks.
@@
II Related work
II-A Scientific language models
The Transformers architecture, along with the growing accessibility of full-text corpora of scientific papers, has heightened the significance of employing NLP techniques for document understanding and knowledge extraction [2]. Similar to general domains, transfer learning through pretraining models can leverage information learned in a self-supervised fashion to aid in downstream scientific NLP tasks where datasets are scarce, leading to the introduction of numerous scientific language models (SciLMs) for a number of fields, including biomedicine [6, 4, 8, 14, 15, 16], chemistry [17], material science [9], and social science [18].
TABLE I: Different domain-specific language models for scientific text.
@@ II-A Scientific language models
Table I summarizes some notable SciLMs that employ the BERT architecture. BioBERT [4] was resulted from continual pretraining of the original BERT model on PubMed abstracts and showed better performances on biomedical NLP tasks. SciBERT [6] was pretrained from scratch on the full-text of 1.1M biomedical and computer science papers in the Semantic Scholar corpus [19]. Evaluations on multiple tasks showed over-performances than the original BERT and comparable performances as BioBERT on biomedical tasks. BiomedBERT [7] was also a pretrained model from scratch on 3.2B tokens in 14M PubMed abstracts. The work highlights the necessity of pretraining from scratch, as it yields a domain specific vocabulary that is useful for downstream tasks. In their case, the vocabulary contains more biomedical terms like “acetyltransferase”, in contrast to the vocabulary in BioBERT that consists of terms common to the general domain. OAG-BERT [5] was trained on the AMiner and PubMed corpora. It is based on the BERT architecture but fine-tuned for academic graph related tasks such as author name disambiguation, paper recommendation, and citation prediction. ScholarBERT [20] was pretrained from scratch on perhaps the largest corpus (221B tokens in 75M journal articles provided by public.resource.org). One observation from this work is that performances in downstream scientific tasks saturate when increasing training data, model size, or training time.
@@
In addition to scientific text, other modalities of scientific data have been used for pretraining language models, including math equations [21], chemicals, proteins, etc. Notably, SMILES strings of chemicals have been utilized to pretrain BERT-like models, yielding several chemical language models [22, 23]. Similarly, protein language models are trained using protein sequences [24].
II-B Temporality of language models
@@
As we shall pretrain SciLMs on temporally organized corpora, our work is broadly related to training models adaptive to diverse domains (e.g., [25], [26]) and specifically related to the temporal aspect of language models. Some studies found that language models implicitly encode various external knowledge including space and time [27]. However, an increasing number of studies reported the temporal misalignment phenomenon, which means that models trained on text from one period can have performance degradation when tested on text from another. For example, [10] showed the effect of temporal drift on NER tasks in temporally-diverse tweets. Likewise, [11] quantified performance degradation due to temporal misalignment across domains and tasks. These studies corroborate the importance of customizing models to different time periods, which is the practice we adopt in this work. To mitigate the effect of temporal misalignment, existing studies suggested a few methods, including continuous pretraining [11, 12], adding year flags to training instances [28], and discarding outdated facts [29]. However, these methods are proposed to improve performances in downstream tasks, but it turned out that the improvement is limited [11].
@@
Another strategy aims at achieving a desired outcome by directly editing a model itself, which involves altering a pretrained model in the weight space to enhance model performance in specific tasks [30, 13]. Current approaches to model editing focused on task arithmetic [30] and weight interpolation [31, 32]. In the context of addressing temporal misalignment, model editing through time vector as a special case of task vector was used to obtain an updated pretrained model for a specific period [13]. However, a critical limitation of time vector based model editing is that the language model used in the first place was pretrained on a time discorded corpus, which confounds the derivation of time vectors. Moreover, using such language model in the scientific domain may be especially problematic, as the corpus size grows drastically due to the exponential growth of science publishing, which means that later corpora would over-shadow earlier ones in determining model weights. By contrast, instead of using model editing, we obtain an updated model through continual training on a series of corpora organized temporally.
@@
In this regard, our pretraining strategy can be considered as a special case of domain adaption [3], where each corpus in a separate period is treated as a “domain”. A natural consequence of domain adaption that was studied is representation forgetting: As a language model adapts to new domains, it forgets some knowledge learned from previous ones, leading to performance degradation [33]. However, in our work, we find that representation forgetting is highly dependent on tasks: We identify two rather similar tasks where for one of them, representation forgetting does manifest, but for the other the opposite of forgetting is observed.
II-C Computational analysis of evolution of science
@@
Distinguishing from the majority of existing SciLMs that concern about downstream NLP tasks, our motivation of pretraining a series of SciLMs is to utilize them to understand quantitatively how science evolves over time through the lens of scientific text. Thus, our work is related to inquiry into the evolution of science, using diverse types of data including full-text and metadata like authorship and citations. [34] proposed dynamic topic models and applied them to analyze the temporal evolution of topics in papers published in the journal Science. Analyzing authors’ transitions between different topics over time, [35] identified distinctive stages in the history of computational linguistics. [36] developed a classifier to assign rhetorical functions (e.g., background, method, conclusion, etc.) to sentences in abstracts and demonstrated the ability to predict the rise and fall of scientific topics by tracking the trajectories of rhetorical functions appeared in different topics over time. [37] developed a classifier to label citation functions (e.g., background, motivation, etc.) and applied it to NLP papers to reveal the evolution of this field.
III Developing AnnualBERT
III-A Preparing corpus
@@
(a) Yearly number of papers by category
(b) Yearly number of sentences and tokens after cleaning
Figure 2: Summary statistics of our arXiv corpus.
@@
To pretrain our models, we first need to prepare a full-text corpus extracted from the downloaded files, which are overwhelmingly and PDF files. For files, we use the LaTeXML tool [38] to convert them into HTML files, which facilitates our followup processing, as there are myriad author-defined commands and symbols in the files. We then remove contents like figures, tables, and equations as well as metadata like author information, acknowledgment, and pagination, by striping away HTML elements corresponding to these contents. Next we replace certain markups with particular tokens, e.g., [EQU] for inline equations, [CITE] for citations, etc., which improves the clarity and utility of the text for training. As for PDF documents, we first use PyMuPDF to extract plain text and employ neatext to further clean elements like tables, emojis, and URLs, considering that PDF files often lack a clear document structure that is conducive to cleaning. We then filter out sentences with less than 20 characters and sentences that contain more than 40% of non-essential elements like punctuation, special characters, stopwords, emojis, and URLs. The code for data processing and training can be found in the https://github.com/jd445/TrainAnuualBERT.
@@
After extracting text from and PDF files, we use NLTK [39] for sentence segmentation and filter out sentences with less than three words. This completes our corpus preparation, and Fig. 2b presents the total number of sentences over time.
@@
Next, we build our customized vocabulary. Existing language models use different tokenization methods to build vocabularies. The original BERT, for example, uses WordPiece, whereas RoBERTa uses BPE. Both tokenization methods consider subword as token, which has the advantage in representing unseen words by combining a list of existing tokens in the vocabulary. Subword tokenization, however, is not suitable for our purpose of examining temporal evolution of scientific discourse at the word level. We thus utilize whole words as tokens, an approach scatteredly taken in previous studies [40]. Particularly, we employ spaCy to tokenize all the abstracts from 1990 to 2020 into words and keep those with frequency $\geq 50$, resulting in a vocabulary with about $53,000$ words. We then add some special tokens including [CITE], [EQU], [FIG], [REF], and [SEC] to the vocabulary for ease of markup representation. This tailored approach ensures that our vocabulary is appropriate for interpretations of terminology pertaining to distinct fields of study.
@@
Fig. 2b presents the total number of tokens over time. Fig. 3 presents the Jaccard similarity matrix between the vocabularies of different BERT models (see § II-A for their details). We note that the similarities between BERT models and the other models (except BioBERT) are generally around 20%. BioBERT and BERT_case have similarity of one, since BioBERT adopts the vocabulary of BERT_case. The similarities among SciBERT, BiomedBERT, and ScholarBERT are approximately 30%, which is higher than that of BERT. This reflects differing vocabulary distributions between general and scientific domains. Our AnnualBERT exhibits lower similarities between the other models, which is attributed to the unique tokenization strategy.
Figure 3: Jaccard similarity matrix between vocabularies.
III-B Pretraining AnnualBERT

### v2

@@ Abstract.
We introduce AnnualBERT, a series of language models designed specifically to capture the temporal evolution of scientific text. Deviating from the prevailing paradigms of subword tokenizations and “one model to rule them all”, AnnualBERT adopts whole words as tokens and is composed of a base RoBERTa model pretrained from scratch on the full-text of 1.7 million arXiv papers published until 2008 and a collection of progressively trained models on arXiv papers at an annual basis. We demonstrate the effectiveness of AnnualBERT models by showing that they not only have comparable performances in standard tasks but also achieve state-of-the-art performances on domain-specific NLP tasks as well as link prediction tasks in the arXiv citation network. Our approach enables the pretrained models to not only improve performances on scientific text processing tasks but also to provide insights into the development of scientific discourse over time. The series of the models is available at https://huggingface.co/jd445/AnnualBERTs.
Scientific language models, BERT, Continuous learning, Temporality.
ccs: Computing methodologies Natural language processingccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Data mining
1. Introduction
@@
Natural language processing (NLP) has been experiencing a new paradigm where Transformers-based language models pretrained on a massive text corpus have significantly improved performance in downstream NLP tasks (Vaswani et al., 2017). In our focused scientific domains, both the encoder and decoder components of Transformers have been extensively employed to develop scientific language models pretrained on the full-text of research papers (Ho et al., 2024).
@@
When developing language models for a target domain, previous studies have suggested a number of pretraining methods. The first is continual training, which uses a model pretrained already on a general domain corpus and continues the pretraining process on the target domain corpus (i.e., domain adaption, mixed-domain pretraining) (Gururangan et al., 2020). BioBERT is one example adapted from the original BERT model to the biomedical domain using the PubMed corpus (Lee et al., 2020). The second approach is training from scratch directly on the target domain corpus. Recent development of BERT-like models, including both multi-domain (Liu et al., 2022b; Beltagy et al., 2019) and domain-specific ones (Gu et al., 2021; Naseem et al., 2021; Trewartha et al., 2022), adopts this strategy to better capture the linguistic and semantic features of their respective fields. Training from scratch has been shown to outperform continual training, because it constructs a vocabulary that encompasses more in-domain tokens, leading to more effective learning from the domain corpus. By contrast, continual training retains the original model’s vocabulary and consequently fragments complex, domain-relevant words into less meaningful tokens. For example, the word “chloramphenicol” is treated as one token by BiomedBERT (PubmedBERT) but six tokens by BioBERT (“ch-lor-amp-hen-ico-l”).
@@
Here we propose a novel pretraining strategy that consists of two stages: domain specific pretraining from scratch followed by progressively continual training on corpora within the same domain organized by year (Fig. 1). Our method is designed to, in addition to recognizing the importance of domain specific pretraining, capture the temporal evolution of language use. Temporal change is one of the most salient features in real-world language data: Scientific papers are published in a chronological order, collectively representing the evolving landscape of scientific knowledge during its development; similarly, news articles and social media posts are produced over time, the collection of which shapes the global news agenda. However, existing efforts, regardless of the training methods, largely ignore the temporal aspect of language data and instead are directed toward developing “one model to rule them all”—one model that can excel in various downstream scientific NLP tasks like named entity recognition.
@@
Figure 1. Workflow for developing our AnnualBERT models pretrained on arXiv papers. We first train from scratch a RoBERTa model on papers published until 2008, denoted as $\mathcal{M}_{\text{base}}$, and then for each subsequent year $t$ from 2009, we use papers published in that year for continual training of $\mathcal{M}_{t-1}$ to obtain $\mathcal{M}_{t}$. During the training process, we use whole word token, represented as $\text{WT}_{1}$, $\text{WT}_{2}$, …, $\text{WT}_{n}$.
@@
By comparison, our proposal enables us to develop a series of language models on scientific documents, each of which represents one historical period and the collection of which allows us to understand the evolution of science. At a broader level, our vision is that a language model itself can be regarded as a highly knowledgeable researcher with extensive reading capabilities across all the scientific fields, who updates its inner workings (i.e., model parameters) upon reading all the papers published in a year. By probing how different versions of the model perform on the same task and by mining how model parameters change, we may be able to uncover how science evolves temporally. From another perspective, language models serve to compress documents over a certain period, thereby reflecting various characteristics of language use during that time, such as vocabulary, semantics, and topics. In this regard, our work is related to some recent studies that identify the temporal misalignment phenomenon (Rijhwani and Preoţiuc-Pietro, 2020; Luu et al., 2022) and the remedies to derive customized models to different periods (Luu et al., 2022; Wang et al., 2023; Nylund et al., 2023). However, the language models used in the first place was pretrained on a time discorded corpus, which confounds the derivation.
@@
  • We evaluate AnnualBERT models on a suit of standard NLP tasks in the scientific domain, showing comparable performances to similar models. Moreover, we formulate several domain-specific tasks and demonstrate that AnnualBERT has better performances for these in-domain tasks.
@@
2. Related work
2.1. Scientific language models
The Transformers architecture, along with the growing accessibility of full-text corpora of scientific papers, has heightened the significance of employing NLP techniques for document understanding and knowledge extraction (Ho et al., 2024). Similar to general domains, transfer learning through pretraining models can leverage information learned in a self-supervised fashion to aid in downstream scientific NLP tasks where datasets are scarce, leading to the introduction of numerous scientific language models (SciLMs) for a number of fields, including biomedicine (Beltagy et al., 2019; Lee et al., 2020; Naseem et al., 2021; Kanakarajan et al., 2021; Luo et al., 2022; Bolton et al., 2024), chemistry (Guo et al., 2021), material science (Trewartha et al., 2022), and social science (Shen et al., 2023).
Table 1. Different domain-specific language models for scientific text.
@@ 2.1. Scientific language models
Table 1 summarizes some notable SciLMs that employ the BERT architecture. BioBERT (Lee et al., 2020) was resulted from continual pretraining of the original BERT model on PubMed abstracts and showed better performances on biomedical NLP tasks. SciBERT (Beltagy et al., 2019) was pretrained from scratch on the full-text of 1.1M biomedical and computer science papers in the Semantic Scholar corpus (Lo et al., 2020). Evaluations on multiple tasks showed over-performances than the original BERT and comparable performances as BioBERT on biomedical tasks. BiomedBERT (Gu et al., 2021) was also a pretrained model from scratch on 3.2B tokens in 14M PubMed abstracts. The work highlights the necessity of pretraining from scratch, as it yields a domain specific vocabulary that is useful for downstream tasks. In their case, the vocabulary contains more biomedical terms like “acetyltransferase”, in contrast to the vocabulary in BioBERT that consists of terms common to the general domain. OAG-BERT (Liu et al., 2022b) was trained on the AMiner and PubMed corpora. It is based on the BERT architecture but fine-tuned for academic graph related tasks such as author name disambiguation, paper recommendation, and citation prediction. ScholarBERT (Hong et al., 2023) was pretrained from scratch on perhaps the largest corpus (221B tokens in 75M journal articles provided by public.resource.org). One observation from this work is that performances in downstream scientific tasks saturate when increasing training data, model size, or training time.
@@
In addition to scientific text, other modalities of scientific data have been used for pretraining language models, including math equations (Peng et al., 2021b), chemicals, proteins, etc. Notably, SMILES strings of chemicals have been utilized to pretrain BERT-like models, yielding several chemical language models (Ross et al., 2022; Ahmad et al., 2022). Similarly, protein language models are trained using protein sequences (Elnaggar et al., 2021).
2.2. Temporality of language models
@@
As we shall pretrain SciLMs on temporally organized corpora, our work is broadly related to training models adaptive to diverse domains (e.g., (Huang et al., 2024), (Fan et al., 2024)) and specifically related to the temporal aspect of language models. Some studies found that language models implicitly encode various external knowledge including space and time (Gurnee and Tegmark, 2023). However, an increasing number of studies reported the temporal misalignment phenomenon, which means that models trained on text from one period can have performance degradation when tested on text from another. For example, (Rijhwani and Preoţiuc-Pietro, 2020) showed the effect of temporal drift on NER tasks in temporally-diverse tweets. Likewise, (Luu et al., 2022) quantified performance degradation due to temporal misalignment across domains and tasks. These studies corroborate the importance of customizing models to different time periods, which is the practice we adopt in this work. To mitigate the effect of temporal misalignment, existing studies suggested a few methods, including continuous pretraining (Luu et al., 2022; Wang et al., 2023), adding year flags to training instances (Dhingra et al., 2022), and discarding outdated facts (Zhang and Choi, 2023). However, these methods are proposed to improve performances in downstream tasks, but it turned out that the improvement is limited (Luu et al., 2022).
@@
Another strategy aims at achieving a desired outcome by directly editing a model itself, which involves altering a pretrained model in the weight space to enhance model performance in specific tasks (Ilharco et al., 2023; Nylund et al., 2023). Current approaches to model editing focused on task arithmetic (Ilharco et al., 2023) and weight interpolation (Matena and Raffel, 2022; Wortsman et al., 2022). In the context of addressing temporal misalignment, model editing through time vector as a special case of task vector was used to obtain an updated pretrained model for a specific period (Nylund et al., 2023). However, a critical limitation of time vector based model editing is that the language model used in the first place was pretrained on a time discorded corpus, which confounds the derivation of time vectors. Moreover, using such language model in the scientific domain may be especially problematic, as the corpus size grows drastically due to the exponential growth of science publishing, which means that later corpora would over-shadow earlier ones in determining model weights. By contrast, instead of using model editing, we obtain an updated model through continual training on a series of corpora organized temporally.
@@
In this regard, our pretraining strategy can be considered as a special case of domain adaption (Gururangan et al., 2020), where each corpus in a separate period is treated as a “domain”. A natural consequence of domain adaption that was studied is representation forgetting: As a language model adapts to new domains, it forgets some knowledge learned from previous ones, leading to performance degradation (Lazaridou et al., 2021). However, in our work, we find that representation forgetting is highly dependent on tasks: We identify two rather similar tasks where for one of them, representation forgetting does manifest, but for the other the opposite of forgetting is observed.
2.3. Computational analysis of evolution of science
@@
Distinguishing from the majority of existing SciLMs that concern about downstream NLP tasks, our motivation of pretraining a series of SciLMs is to utilize them to understand quantitatively how science evolves over time through the lens of scientific text. Thus, our work is related to inquiry into the evolution of science, using diverse types of data including full-text and metadata like authorship and citations. (Blei and Lafferty, 2006) proposed dynamic topic models and applied them to analyze the temporal evolution of topics in papers published in the journal Science. Analyzing authors’ transitions between different topics over time, (Anderson et al., 2012) identified distinctive stages in the history of computational linguistics. (Prabhakaran et al., 2016) developed a classifier to assign rhetorical functions (e.g., background, method, conclusion, etc.) to sentences in abstracts and demonstrated the ability to predict the rise and fall of scientific topics by tracking the trajectories of rhetorical functions appeared in different topics over time. (Jurgens et al., 2018) developed a classifier to label citation functions (e.g., background, motivation, etc.) and applied it to NLP papers to reveal the evolution of this field.
3. Developing AnnualBERT
3.1. Preparing corpus
@@
Figure 2. Summary statistics of our arXiv corpus. (a) Yearly number of papers by category. (b) Yearly number of sentences and tokens after cleaning.
@@
To pretrain our models, we first need to prepare a full-text corpus extracted from the downloaded files, which are overwhelmingly and PDF files. For files, we use the LaTeXML tool (Miller, 2024) to convert them into HTML files, which facilitates our followup processing, as there are myriad author-defined commands and symbols in the files. We then remove contents like figures, tables, and equations as well as metadata like author information, acknowledgment, and pagination, by striping away HTML elements corresponding to these contents. Next we replace certain markups with particular tokens, e.g., [EQU] for inline equations, [CITE] for citations, etc., which improves the clarity and utility of the text for training. As for PDF documents, we first use PyMuPDF to extract plain text and employ neatext to further clean elements like tables, emojis, and URLs, considering that PDF files often lack a clear document structure that is conducive to cleaning. We then filter out sentences with less than 20 characters and sentences that contain more than 40% of non-essential elements like punctuation, special characters, stopwords, emojis, and URLs. The code for data processing and training can be found at https://github.com/jd445/TrainAnuualBERT.
@@
After extracting text from and PDF files, we use NLTK (Loper and Bird, 2002) for sentence segmentation and filter out sentences with fewer than three words. This completes our corpus preparation, and Fig. 2b presents the total number of sentences over time.
Figure 3. Jaccard similarity matrix between vocabularies.
@@
Next, we build our customized vocabulary. Existing language models use different tokenization methods to build vocabularies. The original BERT, for example, uses WordPiece, whereas RoBERTa uses BPE. Both tokenization methods consider subword as token, which has the advantage in representing unseen words by combining a list of existing tokens in the vocabulary. Subword tokenization, however, is not suitable for our purpose of examining temporal evolution of scientific discourse at the word level. Specifically, in § 6.1, we investigate the development of scientific topics by tracking the changing probabilities of a word assigned by different versions of AnnualBERT, a task that is otherwise challenging to accomplish if that word is fragmented into several tokens. We thus utilize whole words as tokens, an approach scatteredly taken in previous studies (Feng et al., 2024). Particularly, we employ spaCy to tokenize all the abstracts from 1990 to 2020 into words and keep those with frequency $\geq 50$, resulting in a vocabulary with about $53,000$ words. We then add some special tokens including [CITE], [EQU], [FIG], [REF], and [SEC] to the vocabulary for ease of markup representation. This tailored approach ensures that our vocabulary is appropriate for interpretations of terminology pertaining to distinct fields of study. Note that words that appear in later corpora bear no influence on earlier models, as the corresponding entries in the embedding layers are kept unchanged as the initial values during model training.
@@
Fig. 2b presents the total number of tokens over time. Fig. 3 presents the Jaccard similarity matrix between the vocabularies of different BERT models (see § 2.1 for their details). We note that the similarities between BERT models and the other models (except BioBERT) are generally around 20%. BioBERT and BERT_case have similarity of one, since BioBERT adopts the vocabulary of BERT_case. The similarities among SciBERT, BiomedBERT, and ScholarBERT are approximately 30%, which is higher than that of BERT. This reflects differing vocabulary distributions between general and scientific domains. Our AnnualBERT exhibits lower similarities between the other models, which is attributed to the unique tokenization strategy.
3.2. Pretraining AnnualBERT